Run plan for 27 September 2002 This should take us through until Monday. Note that Q6 repairs will (almost certainly) take place on Monday... 0. Pulser runs. 1000 events each: - Pulsing chambers - Pulsing postamps - Triggers only (checking for noise) Also make a 3GB file to test the "plain" version of Mofia. 1. Establish system operation - Make sure there are no gas system alarms! - Make sure you can control the M13 beamline. In epics, load up the "m13dev_29p6_all.snap" file, and make sure all the readbacks are close to the set currents. The "fine DAC" controls for B1 and B2 should both be set to 0. - Grant is turning on the chamber HV's and checking their values etc. He and Wayne will do some scope tests while everything else is going on. - Check the postamp thresholds (setting and readback). - The platic degrader is set to position 1 -- ie. the big hole. (I just checked. -Rob.) Check that we have control over the gas degrader (it's on, we can change the mixture, etc). Check that we have signals in the scintillators (ie. they're on). - Check that we can get data with all of our different trigger types: M, MLT, PU, DS. (Take a short run with each. You won't necessarily get the same rate with each trigger, but you should get *something*.) 1a. PACT pedestal runs. See the description linked from "Calibrating the PACT using QOD" linked from the main run page. 2. Efficiency and alignment runs. The efficiency runs are done by varying HV's on groups of planes. Those planes that are not being varied are to be set to the "nominal" values (see the "Nominal Settings" linked from the main run page). Before you start, make sure all the thresholds are set to their "nominal" values. Take 1 Million events (one run) at each setting. The simplest way to do this is to set "Restart: no" on the Midas page. DC EFFICIENCIES With Group A set to 1950 V, take data at Group B HV settings from 2000 V to 1600 V in 50 V steps. Then set Group B to 1950 and do the same with Group A. (We want to scan DC 8 in the first set, you see.) DC Planes in Group A: 1,2,5,6,9,10,13,14,17,18,21,22,25,26,29,30,33,34,37,38,41,42 DC Planes in Group B: 3,4,7,8,11,12,15,16,19,20,23,24,27,28,31,32,35,36,39,40,43,44 PC EFFICIENCIES With Group A set to 2050 V, take data at Group B HV settings from 2200 V to 1600 V in 50 V steps. Then set Group B to 2050 V and do the same with Group A. PC Planes in Group A: 1,2,5,6,9,10 PC Planes in Group B: 3,4,7,8,11,12 We also want a special scan of the target PC's down to 1000 V. Return all other planes (PC's and DC's) to nominal settings, and set the gas degrader to 100% He. Set the target PC's (5, 6, 7, 8) to 1550 V. Wait 90 minutes for the degrader to flush. Then take data at settings from 1550 V to 1000 V at 50 V steps, taking 1 run (1M events) at each setting. When this is done, return the target PC HV's to 2050 V. T0/ALIGNMENT RUNS, etc. Make sure everything is at "nominal" settings. Set M13 to the 120 MeV/c pion spray tune, by loading the following files in epics: m13dev_120_all_slits.snap m13beam_120_all_spray.snap You'll need to go into the area and switch both NMR probes to range 3. You'll also need to put the M13 ABSORBER in (position 1) to get rid of the protons in the beam. Set the trigger to MLT, and take 5 runs. Set the trigger to DS, and take 5 runs. Then tune for 40-50 MeV/c positrons. (You may need to switch NMR probe ranges again.) Leave the slits as they are. Set the trigger to MLT, and take 1 run. Set the trigger to DS, and take 1 run. ---That's it until Monday.